The week began with AR 11041 being the only active region on the Sun, in the SE quadrant. On 20 January it had produced two M-flares while it was at the limb, however, by the start of the planning period (11 UT, 23 Jan) the region was very quiet (few B-flares). Another active region was present in the NW quadrant for the week, but no observations were performed on this.
The week was divided into three OP periods: OP1 was 3 days long, and OP2 and OP3 were 2 days long.
image at 13:13 UT
GOES plot (23-26 Jan)
AR 11041 monitored continuously with FLR001_flare_study, with PRY_slot_context_v3 interspersed every 6 hours or so. I slipped in a FILL001 study at 17:43 to get a spectral atlas:
10:50 FLR001_flare_study (x85)
18:09 FLR001_flare_study (x73)
XRT plan (OP1)
EIT image at 11:48 UT
AR 11041 monitoring with FLR001_flare_study continued until 06:00:
00:04 FLR001_flare_study (x72)
EIT image at 18:24 UT
At 06:46 I slipped in a Eclipse_raster_v2 study to get good spectral coverage of AR 11041. For the SAA-free period I ran cam_artb_lite_v2 for 4 hours, from 09:37.At 18:00 HOP 146 (coronal holes; PI: Ko) was supported at the north pole where the coronal hole was very small.
No observations due to telemetry restrictions.
image at 11:48 UT
GOES plot, 26-28 Jan
A nanoflare study for J. Klimchuk was run from 11:01 to 12:15:
The pointing was set to the east side of the AR at the footpoints of hot loops (moss). A 2nd pointing was performed on 27 Jan (see below). Following this study, single runs of PRY_slot_context_v3 were scheduled every hour before the start of HOP 100.
HOP 100 was supported from 18:12 onwards. SOT takes a large SP map for this, requiring multiple satellite pointings. I requested XRT to run a DEM study, but apparently they're restricted to three filters at the moment. I decided to run a large diagnostic map with EIS. Note that the SOT data were taken between 21:20 and 05:00. The timeline for the observations is
18:12 PRY_slot_contextLITE (x 7)
18:57 HPW017AR_30x400_30s2 (x4, stitching)
20:34 HPW017AR_30x400_30s2 (x4, stitching)
22:11 HPW017AR_30x400_30s2 (x4, stitching)
23:42 HPW017AR_30x400_30s2 (x2, stitching)
00:35 HPW017AR_30x400_30s2 (x1)
01:09 HPW017AR_30x400_30s2 (x2, stitching)
Before each of the HPW017 studies there is a PRY_slot_contextLITE context raster.
The GOES plot shows a number of B-flares between 18:00 and
although these may be from the active region in the NW quadrant.
XRT plan (OP2)
EIT image at 16:25 UT
HOP 100 support was continued with some slot context movies:
01:52 PRY_slot_contextLITE (x 10)
02:42 PRY_slot_contextLITE (x 22)
04:20 PRY_slot_contextLITE (x 13)
05:02 PRY_slot_contextLITE (x 9)
The second pointing (west side of AR) of the Klimchuk nanoflare study was performed between 07:35-08:50:
Three further PRY_slot_context_v3 studies were run at 09:12, 10:15 and 11:15.
No further studies were run except the SYNOP.
A slit alignment study was run at the east limb, which is very quiet:
05:53 SK_DEEP_5x512_SLIT1 (top)
06:04 SK_DEEP_5x512_SLIT1 (bottom)
06:15 SK_DEEP_10x512_SLIT2 (top)
06:20 SK_DEEP_10x512_SLIT2 (bottom)
171 mosaic at 00:16 UT
GOES plot 28-30 Jan
For the period 10:37-15:23 I'm doing an AR outflow study. I first take a detailed raster (PRY_footpoints_HI2, 10:37) at the east footpoints and then a slot context movie (PRY_slot_context_v3, 11:19). The idea is to prove if any flows seen in the context movie actually correspond to blueshifts.
Between 15:30 and 06:00 EIS supports HOP 79 doing a N-S scan.
XRT plan (OP3)
TRACE 171 mosaic at 00:16 UT
HOP 79 is continued to 06:00 today (see above).
For the period 07:10-08:22 I continue the AR outflow study. This time I do the detailed raster (PRY_footpoints_HI2, 07:13) followed by four runs of SK_DEEP_10x512_SLIT2. I position these in pairs such that the first pair covers the top and bottom of the CCD in the proposed outflow region, just to the left of the footpoints. The second pair is positioned 20 arcsec to the left, so further out in the outflow region.
The pointing for PRY_footpoints_HI2 looks like it might be good: see the over-plot on a TRACE 171 image.
HOP 81 (south polar coronal hole) is supported between 10:26 and 15:23.
Between 16:00 and 06:05 I do further AR outflow studies. Firstly, for 16:00-19:09 I do a slot context movie (PRY_slot_context_v3). I then do a high cadence, small raster study with GDZ_LT_PLUME_10x180 for the period 19:25-06:00.
The pointing for the GDZ study was not ideal: see this over-plot on a TRACE 171 image.
TRACE 171 mosaic at 00:26 UT
The AR outflow study from 29 January is continued until 06:00 (see above).